# Word-specific tonal realizations in Mandarin

### v1

@@ Abstract
The pitch contours of Mandarin two-character words are generally understood as being shaped by the underlying tones of the constituent single-character words, in interaction with articulatory constraints imposed by factors such as speech rate, co-articulation with adjacent tones, segmental make-up, and predictability. This study shows that tonal realization is also partially determined by words’ meanings. We first show, on the basis of a Taiwan corpus of spontaneous conversations, using the generalized additive regression model, and focusing on the rise-fall tone pattern, that after controlling for effects of speaker and context, word type is a stronger predictor of pitch realization than all the previously established word-form related predictors combined. Importantly, the addition of information about meaning in context improves prediction accuracy even further. We then proceed to show, using computational modeling with context-specific word embeddings, that token-specific pitch contours predict word type with 50% accuracy on held-out data, and that context-sensitive, token-specific embeddings can predict the shape of pitch contours with 30% accuracy. These accuracies, which are an order of magnitude above chance level, suggest that the relation between words’ pitch contours and their meanings are sufficiently strong to be functional for language users. The theoretical implications of these empirical findings are discussed.
@@ 1 Introduction
Xu, (2001) distinguishes two sources of tonal variation: involuntary and voluntary. Involuntary variation arises from articulatory constraints posited to be beyond the speaker’s control. Voluntary variation arises from the speaker’s communicative intentions, such as eliciting a question or creating emphasis, where changes in rhythm, accent placement, and intonation can all give rise to substantial modification of tonal realization; these voluntary constraints are intimately related to the syntactic and pragmatic functions of an utterance (Gårding, , 1987; Shen, , 1989; Shen, 1990a, ; Xu, , 1999; Liu and Xu, , 2005). In addition, the realization of tone is modulated by sociolinguistic and paralinguistic factors such as gender, dialect, and emotion (Fon and Chiang, , 1999; Zhang et al., , 2006).
@@
Involuntary articulatory constraints on tone production arise both from the segmental makeup of syllables and from the processes of connected speech. At the syllable level, vowels, onsets, and syllable structure all contribute to tonal variation (Howie, , 1974; Ho, , 1976; Whalen and Levitt, , 1995; Xu and Xu, , 2003; Fon and Hsu, , 2007). At the utterance level, where a sequence of tones is produced, the realization of a given tone is greatly influenced by its preceding and following tones, leading to tonal co-articulation (Shih, , 1988; Shen, 1990b, ; Xu, , 1997). Co-articulated tones usually deviate from their canonical tonal shapes; in extreme cases the original shapes are no longer preserved (Shen, , 1989; Shih and Kochanski, , 2000) and may be unrecognisable to native speakers (Xu, , 1994). Physically speaking, it requires a certain amount of time to raise or lower pitch (Xu and Sun, , 2002), and therefore tonal realizations are very dependent on whether speakers have sufficient time to realize a given tonal contour. Under the time pressure of fast speech, tonal targets can usually not be fully realized (Tang and Li, , 2020); this leads to significant deviation from the canonical patterns and often results in tonal reduction (Cheng and Xu, , 2015). The accepted descriptions of the tones as ‘level’, ‘rising’, ‘dipping/low’, and ‘falling’ are therefore generalizations across considerable variability in the fine detail of their phonetic realizations.
@@
In the extensive literature on Mandarin tones, their semantic function is straightforwardly simple: different tones distinguish between alternative meanings. For instance, {CJK*}UTF8bsmi就 jiù ‘then’ and {CJK*}UTF8bsmi九 jiǔ ‘nine’ are differentiated by having a falling and a dipping tone, respectively. However, the very same combinations of segments and tone often realize many other different meanings, as exemplified by {CJK*}UTF8bsmi九 jiǔ ‘nine’ and {CJK*}UTF8bsmi酒 jiǔ ‘alcoholic beverage’. The combination of strong phonotactic constraints on syllable structure and a limited number of lexical tones has given rise to widespread homophony and polysemy, often in combination with homography. For instance, {CJK*}UTF8bsmi就 jiù ‘then’, has a wide range of translation equivalents in English, including then, at once, only, already, to approach, to accomplish, to suffer, and to take advantage of (https://www.pleco.com/, s.v.).
@@
The examples in the preceding paragraph are all monosyllabic words, hence written with a single character; however, in the Chinese Lexical Database (Sun et al., , 2018), only 8% of the 48,000 words are monosyllabic. The majority of Mandarin words are disyllabic, written with two characters. [^1] In this study we use a corpus of Taiwan Mandarin spontaneous speech (Fon, , 2004) and take the word labels supplied by the corpus as given. The labeled words include, for example, nouns such as {CJK*}UTF8bsmi學校 xuéxiào ‘school’, verb forms such as {CJK*}UTF8bsmi學到 xuédào ‘learn+resultative’, and negated verbs such as {CJK*}UTF8bsmi不是 búshì ‘not+be’. Since all these forms have two syllables, we refer to them collectively as disyllabic words. The tonal targets of disyllabic words are taken to be subject to the same voluntary and involuntary constraints that govern the realization of tone in monosyllabic words. As a consequence, all disyllabic words sharing, for example, an initial falling tone and a subsequent rising tone are assumed to be realized with the same underlying pitch contour; any differences in how the tones are realized are assumed to be attributable to the involuntary and voluntary processes described above. However, the present study will show that, alongside the known articulatory constraints, there is a previously undocumented close association between the meanings of Mandarin disyllabic words and the realization of their tonal contours.
@@
The results described in the previous paragraph are compatible with a theory of the mental lexicon that postulates a direct connection between the context-specific meaning of a word token and the details of its phonetic realization. Such a theory has been computationally implemented as the Discriminative Lexicon Model, henceforth DLM (Baayen et al., , 2019; Chuang and Baayen, , 2021; Heitmeier et al., 2023c, ). In this model, lexis and morphology are acquired through a process of error-driven discriminative learning that allows for fine-grained alignments between low-level properties of form and low-level properties of meaning, both operationalized as high-dimensional numeric vectors. The model captures relationships between these vectors in two networks: a comprehension network that maps word form onto word meaning, and a production network that maps word meaning onto word form. The model does not store whole word representations of either kind; rather, the model’s memory consists of a set of networks in which the connection weights are continuously recalibrated with each learning event. In the corresponding theory of the mental lexicon, word forms and meanings do not have representations in memory. The forms are ephemeral auditory or visual experiences, which dynamically generate corresponding, equally ephemeral, meaning representations. Conversely, a meaning conceptualized by a speaker at a given point in time is dynamically transformed into ephemeral representations driving articulation. In other words, the DLM posits a lexicon in which lexical items are neither static nor discrete. Rather, the lexicon is taken to consist of a series of dynamic, modality specific neural networks (Baayen et al., , 2019) which are constantly fine-tuned, by adjusting connection weights, in order to optimize word comprehension and production.
@@
At this point, the question arises of how to understand the linguistic term ‘word’. In this study, we define word token as a pairing of a specific form (which can be represented mathematically as a high-dimensional numeric vector) with a specific meaning (which can also be represented mathematically as a high-dimensional numeric vector). We define word type (or simply, word) as a set of word tokens that have the property of having both similar forms and similar meanings. For instance, the set of phonetic realizations of {CJK*}UTF8bsmi酒 jiǔ and their corresponding context-specific meanings (‘wine, liquor, spirits, alcoholic beverage’, https://www.pleco.com/, s.v.) jointly constitute the tokens of the word type {CJK*}UTF8bsmi酒. [^2] Note that, in this case, the Chinese orthographic character can be used to represent the word, since {CJK*}UTF8bsmi酒 is not homographic: all its meanings cluster around the concept of ‘alcoholic beverage’. This working definition of ‘word’ seeks to do justice to the fact that no two tokens of the same word, as produced by humans, are ever completely identical in form. It also seeks to do justice to the insights from distributional semantics that what a word means varies with its context (Firth, , 1968; Harris, , 1954; Landauer and Dumais, , 1997; Elman, , 2009). Thus, in the framework of the DLM, words are sets of input-output pairs on which the production and comprehension networks are trained, that leave ‘traces’ in the connection weights of these networks, but that are themselves not stored as independent entities.
If the lexicon consists of networks of connection weights, then it is possible that properties of these networks can account for the variation in fine phonetic detail described above. This hypothesis is supported by a series of studies that have used the DLM to model such variation. For example, Gahl and Baayen, (2024) used the DLM to model the different spoken word durations of English homophones, which turn out to be partly determined by how well the spoken form of a homophone token can be predicted from the meaning of that token. Using the same framework, Saito et al., (2023) showed that the strength of connection between features of form and meaning is predictive of the extent to which the tongue is lowered during the articulation of the German [a] vowel, as registered using electromagnetic articulography. For Mandarin, Chuang et al., (2023) were able to predict the duration of words in spontaneous speech in a similar way; as found for other languages, a Mandarin word’s duration tends to be longer if its form and meaning are well aligned, although this pattern appears to be restricted to shorter, semantically transparent words. [^3] In this context, ‘shorter’ words are those whose canonical spoken forms have fewer segments.
@@
Both the theoretical assumptions of the DLM and the empirical studies of English homophoneous words and affixes by Gahl and Baayen, (2024) and Plag et al., (2017) respectively, suggest the possibility that Mandarin homophones also differ systematically in phonetic detail, i.e., that their segments and/or tones are realized slightly differently according to the intended meaning. This could apply to homographic pairs such as {CJK*}UTF8bsmi大家 dàjiā ‘everyone’ and ‘art master’, as well as to non-homographic pairs such as {CJK*}UTF8bsmi樹木 shùmù ‘tree’ and {CJK*}UTF8bsmi數目 shùmù ‘number’. In other words, it is possible that the realizations of canonical tones are determined not only by the involuntary and voluntary constraints previously described, but also by the context-specific meanings of the word tokens on which they are realized. If this is correct, then conversely it is not only the four canonical pitch contours that help to distinguish between alternative meanings, but also the finer details of their phonetic realization. This brings us to the central hypothesis of our study:
@@
  • The unique pitch contour of each spoken Mandarin word token is determined in part by the specific meaning of that token. A speaker can learn to produce meaning-specific pitch contours, and these meaning-specific contours are functional for the listener.
The following four predictions about Mandarin words follow from this hypothesis:
@@
    Variation in the tonal realization of a word cannot be reduced to the segment-related constraints on articulation previously described in the literature.
@@
    Information about a word’s meaning in context will improve prediction of its tonal realization.
@@
    Given a pitch contour, the meaning of its carrier token can be predicted above chance level, assuming the listener has previous experience of that word type.
@@
    Assuming they have previous experience of a given word type, a speaker can produce an appropriate pitch contour for a meaning they want to convey with that word.
In this paper we explore these predictions for disyllabic words with the canonical tone specification of a rising tone (T2) followed by a falling tone (T4), henceforth RF. The disyllabic word is a natural choice for our study since Mandarin vocabulary is composed mostly of disyllabic words (Huang et al., , 2010; Wu et al., , 2023). We decided to focus on the RF pattern, because it is the heterogeneous tonal combination with the highest number of word types and tokens in the speech corpus we used, namely the Taiwan Mandarin spontaneous speech corpus (Fon, , 2004). We wanted to investigate a heterogeneous tonal combination rather than a homogeneous one to ensure that the results obtained are not specific to a given tone.
The remainder of the paper proceeds as follows. Section 2 addresses the first two predictions listed above. It describes how we used generalized additive modeling to analyze the pitch contours of RF words extracted from the above-mentioned corpus of spontaneous speech. We discuss our modeling strategy, and present the results of an analysis based on word type and one enhanced with word sense. These results are then triangulated with a separate analysis using a Random Forest algorithm. Section 3 addresses the third and fourth predictions. It describes how we used computational modeling with the DLM to demonstrate that meaning-specific pitch contours have the potential to facilitate comprehension and to be produced in response to intended meaning. Section 4 is a discussion of the implications of our results.

### v2

@@ Abstract
The pitch contours of Mandarin two-character words are generally understood as being shaped by the underlying tones of the constituent single-character words, in interaction with articulatory constraints imposed by factors such as speech rate, co-articulation with adjacent tones, segmental make-up, and predictability. This study shows that tonal realization is also partially determined by words’ meanings. We first show, on the basis of a corpus of Taiwan Mandarin spontaneous conversations, using a generalized additive regression model, and focusing on the rise-fall tone pattern, that after controlling for effects of speaker and context, word type is a stronger predictor of tonal realization than all the previously established word-form related predictors combined. Importantly, the addition of information about meaning in context improves prediction accuracy even further. We then proceed to show, using computational modeling with context-specific word embeddings, that token-specific pitch contours predict word type with 50% accuracy on held-out data, and that context-sensitive, token-specific embeddings can predict the shape of pitch contours with 40% accuracy. These accuracies, which are an order of magnitude above chance level, suggest that the relation between words’ pitch contours and their meanings are sufficiently strong to be potentially functional for language users. The theoretical implications of these empirical findings are discussed.
@@ 1 Introduction
Not only do tokens of the same word vary substantially in the realization of their segments but, in tonal languages, they also vary substantially in the realization of their lexical tone. Xu, (2001) distinguishes two sources of tonal variation: voluntary and involuntary. Voluntary variation arises from the speaker’s communicative intentions, such as eliciting a question or creating emphasis, where changes in rhythm, accent placement, and intonation can all give rise to substantial modification of tonal realization; these voluntary constraints are intimately related to the syntactic and pragmatic functions of an utterance (Gårding,, 1987; Shen,, 1989; Shen, 1990a, ; Xu,, 1999; Liu and Xu,, 2005). In addition, the realization of tone is modulated by sociolinguistic and paralinguistic factors such as gender, dialect, and emotion (Fon and Chiang,, 1999; Zhang et al.,, 2006).
@@
Involuntary variation in tone production arises from articulatory constraints posited to be beyond the speaker’s control. Some of these constraints are related to the processes of connected speech. At the utterance level, where a sequence of tones is produced, the realization of a given tone is greatly influenced by its preceding and following tones, leading to tonal co-articulation (Shih,, 1988; Shen, 1990b, ; Xu,, 1997). Co-articulated tones usually deviate from their canonical tonal shapes; in extreme cases the original shapes are no longer preserved (Shen,, 1989; Shih and Kochanski,, 2000) and may be unrecognizable to native speakers (Xu,, 1994). Physically speaking, it takes a certain amount of time to raise or lower pitch (Xu and Sun,, 2002), and therefore tonal realizations are highly dependent on whether speakers have sufficient time to realize a given tonal contour. Under the time pressure of fast speech, tonal targets can usually not be fully realized (Tang and Li,, 2020); this leads to significant deviation from the canonical patterns and often results in tonal reduction (Cheng and Xu,, 2015). The accepted descriptions of the tones as ‘level’, ‘rising’, ‘dipping/low’, and ‘falling’ are therefore generalizations across considerable variability in the fine detail of their phonetic realizations.
In addition to the effects of connected speech, articulatory constraints on tone production arise from the segmental makeup of syllables. At the syllable level, vowels, onsets, and syllable structure are all known to contribute to tonal variation (Howie,, 1974; Ho,, 1976; Whalen and Levitt,, 1995; Xu and Xu,, 2003; Fon and Hsu,, 2007). Over and above contextual effects, it is therefore clear that different words with the same canonical tone will differ in the details of their tonal contours, since they differ in their segmental makeup. In previous studies on Mandarin lexical tones using laboratory speech, by-word variation is rarely taken into account, since very often the same words are used across different experimental conditions for maximum control (e.g., Chen,, 2010; Li and Chen,, 2016). Previous studies using corpus data have accommodated by-word variation as a random effect in the relevant statistical models. For example, Wu et al., (2020) worked with random intercepts for word-form. However, this treats the word as a source of noise, where different words exhibit idiosyncrasies that are irrelevant to the predictors of interest. In contrast, other studies have specifically investigated how word-level properties affect tone production. For example, Zhao and Jurafsky, (2009) showed that usage frequency affects F0 realization in Cantonese words with mid-level or mid-rising tones. [^1] Zhao and Jurafsky, (2009) describe Cantonese as having six lexical tones. For Mandarin, Tang and Shaw, (2021) showed that F0 is affected not only by word frequency but also by informativity, defined as a word-level variable. However, to the best of our knowledge, no previous study has specifically addressed the relationship between tonal realization and word meaning. This paper fills that gap.
@@
In the extensive literature on Mandarin tones, their semantic function is straightforwardly simple: different tones distinguish between alternative meanings. For instance, {CJK*}UTF8bsmi就 jiu4 ‘then’ and {CJK*}UTF8bsmi九 jiu3 ‘nine’ are differentiated by having a falling and a dipping tone, respectively. However, the very same combinations of segments and tone often realize many other different meanings, as exemplified by {CJK*}UTF8bsmi九 jiu3 ‘nine’ and {CJK*}UTF8bsmi酒 jiu3 ‘alcoholic beverage’. [^2] Throughout this paper, pinyin transcriptions are followed by numeric notations representing the four tones. The combination of strong phonotactic constraints on syllable structure and a limited number of lexical tones has given rise to widespread homophony and polysemy, often in combination with homography. For instance, {CJK*}UTF8bsmi就 jiu4 ‘then’, has a wide range of translation equivalents in English, including then, at once, only, already, to approach, to accomplish, to suffer, and to take advantage of (https://www.pleco.com/, s.v.).
@@
The examples in the preceding paragraph are all monosyllabic words; however, in the Chinese Lexical Database (Sun et al.,, 2018), only 8% of the 48,000 words are monosyllabic. The majority (70.4%) of Mandarin words are disyllabic, written with two characters. [^3] In this study we use a corpus of Taiwan Mandarin spontaneous speech (cf. Fon,, 2004) and take the word labels supplied by the corpus as given. The labeled words include, for example, nouns such as {CJK*}UTF8bsmi學校 xue2xiao4 ‘school’, verb forms such as {CJK*}UTF8bsmi學到 xue2dao4 ‘learn+resultative’, and negated verbs such as {CJK*}UTF8bsmi不是 bu2shi4 ‘not+be’. Since all these forms have two syllables, we refer to them collectively as disyllabic words. The tonal targets of disyllabic words are taken to be subject to the same voluntary and involuntary constraints that govern the realization of tone in monosyllabic words. As a consequence, all disyllabic words sharing, for example, an initial falling tone and a subsequent rising tone are assumed to have the same underlying pitch contour; any differences in how the tones are realized are assumed to be attributable to the involuntary and voluntary processes described above. However, the present study will show that, alongside the known articulatory constraints, there is a previously undocumented close association between the meanings of Mandarin disyllabic words and the realization of their tonal contours.
@@
The results described in the previous paragraph are compatible with a theory of the mental lexicon that postulates a direct connection between the context-specific meaning of a word token and the details of its phonetic realization. Such a theory has been computationally implemented as the Discriminative Lexicon Model, henceforth DLM (Baayen et al.,, 2019; Chuang and Baayen,, 2021; Heitmeier et al.,, 2025). In this model, lexis and morphology are acquired through a process of error-driven learning that allows for fine-grained alignments between low-level properties of form and low-level properties of meaning, both operationalized as high-dimensional numeric vectors. The model captures relationships between these vectors in two networks: a comprehension network that maps word form onto word meaning, and a production network that maps word meaning onto word form. In other words, given a word form vector, the comprehension network is able to compute a predicted meaning vector, analogous to the comprehension process in which we recognize and make sense of the meaning of a word according to the visual or auditory input. In the same vein, the production network computes a predicted form vector given a semantic vector, analogous to the production process where we produce a form to express an intended word meaning. It has been shown that, across a range of languages, these networks successfully capture an alignment between meaning and fine-grained variation in form, e.g., word duration in Mandarin (Chuang et al.,, 2023), different durations of homophones in English (Gahl and Baayen,, 2024), and the degree of tongue lowering in articulation of the [a] vowel in German (Saito et al.,, 2023).
Despite being able to predict both word forms and word meanings, the DLM model does not store whole word representations of either kind; rather, the model’s memory consists of the connection weights in the networks, which are continuously re-calibrated with each learning event. Likewise, in the corresponding theory of the mental lexicon, word forms and meanings do not have representations in memory. The forms are ephemeral auditory or visual experiences, which dynamically generate corresponding, equally ephemeral, meaning representations. Conversely, a meaning conceptualized by a speaker at a given point in time is dynamically transformed into ephemeral representations driving articulation. In other words, the DLM posits a lexicon in which lexical items are neither static nor discrete. Rather, the lexicon is taken to consist of a series of dynamic, modality specific neural networks (Baayen et al.,, 2019) which are constantly fine-tuned, by adjusting connection weights, in order to optimize word comprehension and production. For further technical details of the implementation of the DLM mappings, please refer to the supplementary material.
@@
At this point, the question arises of how to understand the linguistic term ‘word’. In this study, we define word token as a pairing of a specific form with a specific meaning. We define word type (or simply, word) as a set of word tokens that have the property of having both similar forms and similar meanings. For instance, the set of phonetic realizations of {CJK*}UTF8bsmi酒 jiu3 and their corresponding context-specific meanings (‘wine, liquor, spirits, alcoholic beverage’, https://www.pleco.com/, s.v.) jointly constitute the tokens of the word type {CJK*}UTF8bsmi酒. [^4] Note that, in this case, the Chinese orthographic character can be used to represent the word, since {CJK*}UTF8bsmi酒 is not homographic: all its meanings cluster around the concept of ‘alcoholic beverage’. This working definition of ‘word’ seeks to do justice to the fact that no two tokens of the same word, as produced by humans, are ever completely identical in form. It also seeks to do justice to the insights from distributional semantics that what a word means varies with its context (Firth,, 1968; Harris,, 1954; Landauer and Dumais,, 1997; Elman,, 2009). Thus, in the framework of the DLM, words are sets of input-output pairs on which the production and comprehension networks are trained, that leave ‘traces’ in the connection weights of these networks, but that are themselves not stored as independent entities.
As may be apparent from the preceding paragraph, our theoretical position apropos meaning is one of contextualism. We assume that utterances rather than sentences are the domain of propositional content, and that the meaning of an utterance depends on the context in which it is produced. Our notion of context includes not only the content, genre and style of the surrounding text but also extralinguistic factors such as who is speaking and who they are addressing. Hence, our view of meaning does not draw a neat boundary between semantics and pragmatics.
@@
Both the theoretical assumptions of the DLM and the empirical studies of English homophonous words and affixes by Gahl and Baayen, (2024) and Plag et al., (2017) respectively, suggest the possibility that Mandarin homophones also differ systematically in phonetic detail, i.e., that their segments and/or tones are realized slightly differently according to the intended meaning. This could apply to homographic pairs such as {CJK*}UTF8bsmi大家 da4jia1 ‘everyone’ and ‘art master’, as well as to non-homographic pairs such as {CJK*}UTF8bsmi樹木 shu4mu4 ‘tree’ and {CJK*}UTF8bsmi數目 shu4mu4 ‘number’. In other words, it is possible that the realizations of canonical tones are determined not only by the involuntary and voluntary constraints previously described, but also by the context-specific meanings of the word tokens on which they are realized. If this is correct, then conversely it is not only the four canonical pitch contours that help to distinguish between alternative meanings, but also the finer details of their phonetic realization. This brings us to the central hypothesis of our study:
@@
  • The unique pitch contour of each spoken Mandarin word token is determined in part by the specific meaning of that token. These meaning-specific contours are in principle learnable and hence potentially used by speakers.
In what follows, we provide evidence for four more specific predictions that we derive from this hypothesis:
@@
    Word type will be a stronger predictor of tonal realization than all the previously established word-form related predictors combined. This prediction follows from the hypothesis because word type includes information about meaning in addition to information about form.
@@
    Information about a word’s meaning in context, i.e. its sense, will improve prediction of its tonal realization, compared with prediction based on word type alone. This prediction follows from the hypothesis because a word’s senses include all the information encoded in the word, plus additional more fine-grained semantic information. It therefore allows us to tease apart effects of meaning from possible effects of form.
@@
    Given a pitch contour, the meaning of its carrier token can be predicted above chance level by a simple computational model with previous experience of that word type. This prediction follows from the hypothesis that meaning-specific contours are in principle learnable.
@@
    Assuming it has previous experience of the relevant word type, a simple computational model can produce an appropriate pitch contour for a given meaning. This also follows from the hypothesis that meaning-specific contours are in principle learnable.
In this paper we explore these predictions for disyllabic words with the canonical tone specification of a rising tone (T2) followed by a falling tone (T4), henceforth RF. The disyllabic word is a natural choice for our study since Mandarin vocabulary is composed mostly of disyllabic words (Huang et al.,, 2010; Wu et al.,, 2023). We decided to focus on the RF pattern, because it is the heterogeneous tonal combination with the highest number of word types and tokens in the speech corpus we used. We wanted to investigate a heterogeneous tonal combination rather than a homogeneous one to ensure that the results obtained are not specific to a given tone.
The remainder of the paper proceeds as follows. Section 2 addresses the first two predictions listed above. It describes how we used generalized additive modeling to analyze the pitch contours of RF words extracted from the above-mentioned corpus of spontaneous speech. We discuss our modeling strategy, and present the results of an analysis based on word type and one enhanced with word sense. Section 3 addresses the third and fourth predictions. It describes how we used computational modeling with the DLM to demonstrate that meaning-specific pitch contours have the potential to facilitate comprehension and to be produced in response to intended meaning. Section 4 is a discussion of the implications of our results.